# Efficient Exact Algorithms for Minimum Covering of Orthogonal Polygons with Squares

### v1

@@ 3 Structural and Geometric Results
In this section, we prove some structural and geometric results of a minimum square covering of an orthogonal polygon $P$, as well as a set of rec-packs which is a minimum covering of $P$.
@@ 3.1 Structure of Minimum Coverings
Due to Figure 3 we assume that we are looking for a minimum covering where every square is a maximal square. First, let us define how a maximal square can be obtained from a convex vertex of the given orthogonal polygon $P$.
Definition 3.1.
Given an orthogonal polygon $P$ and a point $v$, the following process is called growing of a square with one corner fixed at $v$:
  • start with $S_{1}$ being the $1\times 1$ valid square covering $v$ (and has $v$ as one of its corners).
  • for $t$ in $2,3,\ldots$, from $S_{t-1}$, which is a $(t-1)\times(t-1)$ square with $v$ as one of its corners, construct $T$ to be a $t\times t$ square by keeping one of its vertex fixed at $v$ and changing both the $x$- and $y$-coordinates of the diagonally opposite vertex in $S_{t-1}$ by $1$ to obtain the corresponding diagonally opposite vertex in $T$. If $T$ is a valid square, set $S_{t}\leftarrow T$; else terminate the process and set $S=S_{t-1}$.
$S$ is said to be the square grown from the point $v$.
We now discuss the following result.
@@
Given an orthogonal polygon $P$, the maximal square covering a convex vertex is unique. Further, for a given convex vertex $v$, the unique maximal square can be found in $\mathcal{O}(n)$.
Proof 3.3.
Given a convex vertex $v$ in an orthogonal polygon $P$, let $S$ be the vertex grown from $v$. As all other squares covering $v$ strictly lie inside $S$, $S$ must be the unique maximal square covering $v$.
Let $\mathcal{R}$ denote the region defined by the interior of the $90^{\circ}$ angle between the rays formed by extending the two edges adjacent to $v$ (i.e. a quadrant). To find $S$ in $\mathcal{O}(n)$ time, we can iterate over all polygon edges $(v_{i},v_{i+1})$, and check the following:
@@
  • if some portion of $(v_{i},v_{i+1})$ lies inside $\mathcal{R}$, then find in $\mathcal{O}(1)$, the largest square in $\mathcal{R}$, with one vertex at $v$ and the strict interior of the square does not contain any portion of $(v_{i},v_{i+1})$. This can be done in $\mathcal{O}(1)$, just by comparing the coordinates of $v_{i}$, $v_{i+1}$ and $v$.
@@
We now define a maximal corner square of a vertex as follows.
@@
If $S$ is a maximal square of an orthogonal polygon $P$, then either both the vertical sides of $S$ overlap fully or partially with some polygon edges in $P$ or both the horizontal sides of $S$ overlap fully or partially with some polygon edges of $P$.
Proof 3.7.
Assume the contrary. Suppose for some maximal square $S$ there is one horizontal edge and one vertical edge which does not overlap fully or partially with any polygon edges of $P$. Further, we can assume that these are the top and the right edges of $S$ (we can rotate $P$ along with $S$ appropriately if this is not the case). Then we can further grow $S$ fixing its bottom-left corner, which contradicts that $S$ is a maximal square.
@@ 3.2 Simplicial Nodes in the Associated Graph and Partial Solutions
Given an orthogonal polygon $P$, a set of rec-packs $\mathcal{R}$ is a partial solution for OPCS if there is a minimum covering set of rec-packs $\mathcal{R}^{\prime}$ with $\mathcal{R}\subseteq\mathcal{R}^{\prime}$.
Now consider a partial solution $\mathcal{R}$. We define $G^{\mathcal{R}}(P)$ as follows.
@@
Let $G^{\mathcal{R}}(P)$ denote the induced subgraph of the associated graph $\mathcal{G}(P)$ consisting of nodes which correspond to blocks in $P$ which are not covered by any rec-pack in $\mathcal{R}$.
From Proposition 2.16 and Proposition 2.14, we get that $G^{\mathcal{R}}(P)$ is chordal, and from Proposition 2.15 we get that $G^{\mathcal{R}}(P)$ has a simplicial node $p$ if $G^{\mathcal{R}}(P)$ is non-empty. Let $A$ denote the union of the block $p$ and its neighbouring blocks in $G^{\mathcal{R}}(P)$. By Definition 2.13, $A$ induces a clique in $G^{\mathcal{R}}(P)$ and hence in $G(P)$. Therefore, by Proposition 1.2, there exists a maximal square $S_{A}$ that covers all blocks in $A$. Next, we define unambiguous squares given a partial solution $\mathcal{R}$.
@@
Let $\mathcal{R}$ be a partial solution for OPCS in an orthogonal polygon $P$. We call a maximal square $S$ to be an unambiguous square given $\mathcal{R}$, if there is simplicial node $p$ in $G^{\mathcal{R}}(P)$ such that $S$ covers $p$ and all its neighbours in $G^{\mathcal{R}}(P)$.
@@
For a given simplicial node $p$ in $G^{\mathcal{R}}(P)$, there can be multiple maximal squares that cover $p$ and all its neighbours in $G^{\mathcal{R}}(P)$.
We now justify the naming of unambiguous squares.
@@
If $\mathcal{R}$ is a partial solution for OPCS in an orthogonal polygon $P$ and $S$ is an unambiguous square given $\mathcal{R}$, then $\mathcal{R}\cup\{S\}$ is also a partial solution.
Proof 3.13.
Let $p$ be a simplicial node in $G^{\mathcal{R}}(P)$ such that $S$ covers $p$ and its neighbours in $G^{\mathcal{R}}(P)$. Since $\mathcal{R}$ is a partial solution, there exists a minimum cover $\mathcal{R}^{\prime}$ covering $P$ with $\mathcal{R}\subseteq\mathcal{R}^{\prime}$. This implies $\mathcal{E}(\mathcal{R})\subseteq\mathcal{E}(\mathcal{R}^{\prime})$ from Definition 2.9. Let $S^{\prime}\in\mathcal{E}(\mathcal{R}^{\prime})$ be a maximal square covering $p$. We have $S^{\prime}\notin\mathcal{E}(\mathcal{R})$ as $p$ is in $G^{\mathcal{R}}(P)$. This implies that ()()$\mathcal{E}(\mathcal{R})\subseteq\mathcal{E}(\mathcal{R}^{\prime})\setminus\{S^{\prime}\}$. However, since $S^{\prime}$ covers an induced clique of $G^{\mathcal{R}}(P)$ including $p$ (Proposition 1.2) and possibly some blocks already covered by squares in $\mathcal{E}(\mathcal{R})$, this means that $S^{\prime}$ covers $p$ and a subset of its neighbours in $G^{\mathcal{R}}(P)$ along with some blocks already covered by squares in $\mathcal{E}(\mathcal{R})$. On the other hand, recall that $S$ covers $p$ and all its neighbours in $G^{\mathcal{R}}(P)$. This means ()$\mathcal{E}(\mathcal{R})\cup\{S^{\prime}\}$ covers a subset of points covered by ()$\mathcal{E}(\mathcal{R})\cup\{S\}$. This means (()$(\mathcal{E}(\mathcal{R}^{\prime})\setminus\{S^{\prime}\})\cup\{S\}$ is also a minimum cardinality set of squares covering the entire polygon $P$. Let $\mathcal{Z}=(\mathcal{E}(\mathcal{R}^{\prime})\setminus\{S^{\prime}\})\cup\{S\}$.
We rewrite $\mathcal{Z}=(\mathcal{E}(\mathcal{R}^{\prime})\setminus\{S^{\prime}\})\cup\{S\}=(\mathcal{E}(\mathcal{R})\cup\mathcal{T})\cup\{S\}$, where $\mathcal{T}=(\mathcal{E}(\mathcal{R}^{\prime})\setminus\{S^{\prime}\})\setminus\mathcal{E}(\mathcal{R})$. Consider $\mathcal{Z}_{1}=\mathcal{R}\cup\mathcal{T}\cup\{S\}$ to be a set of rec-packs (where $\mathcal{T}$ and $\{S\}$ are considered to be sets of trivial rec-packs). Since $\mathcal{E}(\mathcal{Z}_{1})=\mathcal{E}(\mathcal{R}\cup\mathcal{T}\cup\{S\})=\mathcal{E}(\mathcal{R})\cup\mathcal{T}\cup\{S\}=\mathcal{Z}$ is a minimum cardinality set of squares covering $P$, $\mathcal{Z}_{1}$ is a minimum covering. Moreover, as $\mathcal{R}\cup\{S\}\subseteq\mathcal{Z}_{1}$, $\mathcal{R}\cup\{S\}$ is a partial solution. Hence, we are done.
Now, we prove that we can always find an unambiguous square which follows a particular structure.
@@
Let $\mathcal{R}$ be a partial solution of OPCS in an orthogonal polygon $P$ such that $\mathcal{E}(\mathcal{R})$ does not completely cover $P$. We define $C_{x}$ and $C_{y}$ as a set of lattice points as follows.
@@
$y\text{ is a y-coordinate of a vertex of }P\text{ or a corner ofa rec-pack in }\mathcal{R}\}$
@@
$y\text{ is a y-coordinate of a vertex of }P\}$
Then there exists an unambiguous square $S$ given $\mathcal{R}$ which has one corner in $C_{x}\cup C_{y}$.
Proof 3.15.
Let $p$ be a simplicial node in $G^{\mathcal{R}}(P)$ with the largest number of neighbours, and such that among the topmost such simplicial nodes it is the leftmost. Let $A$ denote the following set of nodes (blocks in $P$): $p$ and its neighbours in $G^{\mathcal{R}}(P)$. Let $S_{0}$ be a maximal square that covers all blocks in $A$, such that the position of the top-left corner of $S_{0}$ is leftmost among the topmost of all such squares covering $A$. By Lemma 3.6, we know that either the horizontal sides of $S_{0}$ overlaps (completely or partially) with two horizontal polygon edges, or the vertical sides of $S_{0}$ overlaps (completely or partially) with two vertical polygon edges.
Assume that the horizontal sides of $S_{0}$ overlaps (completely or partially) with two horizontal polygon edges. We consider two cases.
@@
Since $p$ is a simplicial node with the largest number of neighbours in $G^{\mathcal{R}}(P)$, all simplicial nodes (in fact all nodes) in $G^{\mathcal{R}}(P)$ must be isolated. Further, $p$ is the leftmost among the topmost such simplicial nodes. This means that the immediate left block $p^{\prime}$ of $p$ is either covered by a rec-pack in $\mathcal{R}$ or is outside the polygon. Therefore, the left side of $S_{0}$ is either a polygon edge (if $p$ is in the polygon but $p^{\prime}$ is not) or a vertical side of some rec-pack in $\mathcal{R}$ (if $p$ is uncovered, but $p^{\prime}$ is covered), respectively. Therefore, the top-left corner of $S_{0}$ must share the x-coordinate of either a corner of a rec-pack in $\mathcal{R}$ or a polygon vertex. Further, the top-left corner of $S_{0}$ shares a y-coordinate of a polygon vertex as the horizontal sides (top and bottom) of $S_{0}$ overlap with polygon edges. This means that the top-left vertex of $S_{0}$ is in $C_{y}$
@@
We already know that the horizontal sides (top and bottom) of $S_{0}$ overlap with polygon edges. If the left (or right) side of $S_{0}$ shares x-coordinate of some polygon edge or a corner of some rec-pack in $\mathcal{R}$, then the top-left (or top-right) corner of $S_{0}$ is in $C_{y}$; and then we are done.
Now, we assume the contrary. Suppose neither the left or right sides of $S_{0}$ share an x-coordinate with a polygon edge or a corner of some rec-pack in $\mathcal{R}$. In particular, if $S_{l}$ and $S_{r}$ are squares congruent to $S_{0}$ but are shifted one unit to the left and right respectively, then $S_{l}$ and $S_{r}$ are maximal valid squares (Figure 6(a)). They are valid as the left and right sides of $S_{0}$ do not overlap with any polygon edge. They are also maximal as the same polygon edges overlapping horizontal sides of $S_{0}$ also overlap the horizontal sides of $S_{l}$ and $S_{r}$ (this would not be the case if the horizontal polygon edges only overlap horizontal sides of $S_{0}$ at just a point — i.e. a corner of $S_{0}$ overlaps with a vertex of $P$; but then the vertical side of that corner of $S_{0}$ would share the x-coordinate of the vertex of $P$, contradicting our assumption). We further consider three more subcases.
  • Case II(a): All blocks in $S_{0}$ along its right boundary are also covered by some rec-pack in $\mathcal{R}$. This case is impossible as $S_{l}$ covers the same set of uncovered blocks but has a top-left corner which lies even to the left of $S_{0}$ contradicting the definition of $S_{0}$.
  • Case II(b): All blocks in $S_{0}$ along its left boundary are also covered by some rec-pack in $\mathcal{R}$. We keep moving $S_{0}$ horizontally to the right till one of the following happens:
    • (i) it is obstructed by a polygon edge to its right (then the right side would share x-coordinate of a polygon edge)
    • (ii) moving it any further uncovers some previously covered portion (then the left side would be overlapping with the right side of some rec-pack in $\mathcal{R}$, hence sharing the x-coordinate of a corner of some rec-pack in $\mathcal{R}$)
    • (iii) it is about to be no longer overlapping with the horizontal polygon edges of $P$ which overlapped horizontal sides of $S_{0}$ (then there would be a corner of the square that overlaps with the endpoint of the horizontal polygon edge which was about to lose overlap and hence they share the x-coordinate)
    Since the horizontal sides (top and bottom) of $S_{0}$ overlap with polygon edges, in each of the above cases there is a corner of the newly moved square which lies in $\mathcal{C}_{y}$.
@@
    (b) Blocks $p_{l},p_{ll},p_{r},p_{rr}$ given $S_{0}$ and rec-packs in $\mathcal{R}$

### v2

@@ 2 Preliminaries
3 Overview of the Algorithms and the Reduction
3.1 Polynomial-time algorithm for OPCS
The algorithm by Bar-Yehuda and Ben-Chanoch [6] starts with an empty set and keeps adding essential or unambiguous squares one by one, while maintaining the invariant that there is always a minimum covering containing the current set of squares. This algorithm runs in time $\mathcal{O}(n+\text{OPT})$, OPT being the minimum number of squares needed to cover the polygon. Recall that OPT may not be bounded by a polynomial in $n$. Our approach will be the same, except along with unambiguous squares, we will also use rec-packs while building our solution. This is because, our aim is to design an algorithm that runs in time polynomial in the number of vertices $n$ of the polygon.
The basic framework will be to keep adding squares/rec-packs to the existing set of rec-packs, that preserve the invariant: the currently constructed set of rec-packs can always be extended to a minimum covering of $P$. We stop as soon as the set of rec-packs cover the entire polygon $P$; by the invariance property, this must be a minimum covering.
The primary challenge for designing an algorithm that runs in time polynomial in the number of vertices, lies in finding unambiguous squares if they exist, and even figuring out what happens if there are no unambiguous squares. Firstly, we prove that there will always be at least one unambiguous square. This comes from the fact that there will always be a simplicial node in any induced subgraph of the chordal graph $G(P)$. However, finding an unambiguous square is still not trivial to achieve in time polynomial in $n$.
We break down the task of finding unambiguous squares into two subtasks: (i) obtaining a polynomial-sized set of squares which contain at least one unambiguous square, and (ii) detecting which of them is unambiguous.
Using this framework directly would still give us an output-sensitive algorithm, since we are enumerating each square one by one. However, this can be sped up by including rec-packs to the partial solution. Suppose at an instant, the algorithm finds an unambiguous square $S$ and adds it to the current set of rec-packs. At such a stage, let $\mathscr{R}$ be the set of rec-packs currently constructed by our algorithm, which can be extended to a minimum covering $\mathscr{R}^{\star}$ of $P$. We ensure that the algorithm further detects if $S$ is in some non-trivial rec-pack $R$ in $\mathscr{R}^{\star}$. In such a case, we replace the square $S$ by the rec-pack $R$ in our current set of rec-packs. Since $R$ may contain multiple new squares, it is possible that this step includes multiple squares to our solution in a single iterations, and this might result in a reduction in the number of iterations. In fact, we show that $\mathcal{O}(n^{2})$ iterations are enough, making the running time guarantee of the algorithm to be polynomial in $n$, and independent of the output. The exact time complexity turns out to be $\mathcal{O}(n^{10})$; details of this are discussed in Section 5.
3.2 Recursive algorithm using separating squares
Next, we consider the problem $p$-OPCS, where the number of knobs is at most $k$. We design a recursive algorithm for $p$-OPCS which is faster than our previous algorithm if $k$ is small enough.
For our recursive algorithm, we crucially use the structure of separating squares: maximal squares that have a corner at a non-knob convex vertex of $P$, and deletion of which separates the polygon. Let $P$ have $l$ non-knob convex vertices. We find a separating square $S$, the deletion of which separates the polygon into smaller polygons $Q^{\prime}_{1},Q^{\prime}_{2},\ldots,Q^{\prime}_{t^{\prime}}$, $t^{\prime}\leq n$. We now appropriately group $Q^{\prime}_{1},\ldots Q^{\prime}_{t^{\prime}}$ to get $Q_{1},\ldots,Q_{t}$, $t\leq 12$ such that each of the polygons $Q_{1}\cup S,Q_{2}\cup S,\ldots,Q_{t}\cup S$ have at most $k$ knobs, and at most $l-1$ non-knob convex vertices. This allows us to recurse on the polygons $Q_{1}\cup S,Q_{2}\cup S,\ldots,Q_{t}\cup S$. The base case of the recursion is when there are no non-knob convex vertices, and this is solved by the algorithm of Section 5.
The running time analysis of this recursive framework highly relies on the following results:
  • There are $\mathcal{O}(n)$ recursive steps and $\mathcal{O}(n)$ base cases
  • Each recursive step requires $\mathcal{O}(n)$ time
  • The number of vertices for a polygon that appears as a base case is $\mathcal{O}(k)$
This framework therefore gives us a running time guarantee of $\mathcal{O}(n^{2}+n\cdot k^{10})$. This is discussed in detail, in Section 6.
3.3 Reducing OPCSH from Planar 3-CNF
The proof of NP-hardness of OPCSH by Aupperle et. al [4] incorrectly reduces from the problem of the existence for a tautology of a Planar 3-CNF instance (linear-time solvable), instead of the problem of satisfiability of a Planar 3-CNF instance (NP-complete [15]). This is because the gadget used for clauses can be covered by 12 squares if all literals appear as true; otherwise 13 squares are required. Intuitively, when the literals of some clause are such that some evaluate to true and some evaluate to false, the number of squares required to cover the gadget must be less than when all literals are false. This is because the former setting makes the clause evaluate to true, but the latter makes the clause evaluate to false.
We reduce OPCSH from satisfiability of Planar 3-CNF. In our construction, just like in [4] we have variable gadgets, clause gadgets and connector gadgets. Our variable and connector gadgets are exactly the same as those in [4]. However, we introduce a novel construction for a clause gadget. Our construction requires 29 squares to cover a clause gadget, if all literals appear as false; otherwise 28 squares are required. This completes the proof of NP-hardness of OPCSH. This is discussed in detail, in Section 7.
@@ 4 Structural and Geometric Results
We prove some structural and geometric results of a minimum square covering of an orthogonal polygon $P$, and those of a set of rec-packs forming a minimum covering of $P$.
@@ 4.1 Structure of minimum coverings
Due to Section 2 we direct our focus to minimum coverings where every square is a maximal square. First, we define how a maximal square can be obtained from a convex vertex of the given orthogonal polygon $P$.
@@
The maximal square of an orthogonal polygon $P$, covering a convex vertex $v$ is unique and can be found in $\mathcal{O}(n)$.
Proof 4.2.
Let $\mathcal{R}$ denote the region defined by the interior of the $90^{\circ}$ angle between the rays formed by extending the two edges adjacent to $v$ (i.e. a quadrant). The largest valid square $S$ in this quadrant with one corner at $v$ is the unique maximal square covering $v$.
To algorithmically find $S$ in $\mathcal{O}(n)$ time, we iterate over all polygon edges $(v_{i},v_{i+1})$, and check the following:
@@
  • if some portion of $(v_{i},v_{i+1})$ lies inside $\mathcal{R}$, the largest square in $\mathcal{R}$, with one vertex at $v$ and the strict interior of the square does not overlap with $(v_{i},v_{i+1})$. This can be done in $\mathcal{O}(1)$, just by comparing the coordinates of $v_{i}$, $v_{i+1}$ and $v$.
@@
We define a maximal corner square of a vertex as follows.
@@
If $S$ is a maximal square of an orthogonal polygon $P$, then either both the vertical sides of $S$ overlap with some polygon edges in $P$ or both the horizontal sides of $S$ overlap with some polygon edges of $P$.
Proof 4.6.
Assume the contrary. Suppose for some maximal square $S$ there is one horizontal side and one vertical side which does not overlap with any polygon edges of $P$. Further, we can assume that these are the top and the right edges of $S$ (Proposition 2.1). Then we can further grow $S$ fixing its bottom-left corner, which contradicts that $S$ is a maximal square.
@@ 4.2 Simplicial nodes in the associated Graph and partial solutions
Given an orthogonal polygon $P$, a set of rec-packs $\mathscr{R}$ is a partial solution for OPCS if there is a minimum covering set of rec-packs $\mathscr{R}^{\prime}$ with $\mathscr{R}\subseteq\mathscr{R}^{\prime}$.
Now consider a partial solution $\mathscr{R}$. We define $G^{\mathscr{R}}(P)$ as follows.
@@
Let $G^{\mathscr{R}}(P)$ denote the induced subgraph of the associated graph $G(P)$ consisting of nodes corresponding to blocks in $P$ which are not covered by rec-packs in $\mathscr{R}$.
Proposition 2.17 and Proposition 2.15 imply that $G^{\mathscr{R}}(P)$ is chordal, and Proposition 2.16 implies that $G^{\mathscr{R}}(P)$ has a simplicial node $p$ if $G^{\mathscr{R}}(P)$ is non-empty. Note that nodes in $G^{\mathscr{R}}(P)$ or $G(P)$ are blocks in $P$. Let $A$ denote the union of the block $p$ and its neighbouring blocks in $G^{\mathscr{R}}(P)$. $A$ induces a clique in $G^{\mathscr{R}}(P)$ and hence in $G(P)$ (Definition 2.14). Therefore, there exists a maximal square $S_{A}$ that covers all blocks in $A$ (Proposition 1.2). With this, we define unambiguous squares given a partial solution $\mathscr{R}$.
@@
Let $\mathscr{R}$ be a partial solution for an orthogonal polygon $P$. We call a maximal square $S$ to be an unambiguous square given $\mathscr{R}$, if there is simplicial node $p$ in $G^{\mathscr{R}}(P)$ such that $S$ covers $p$ and all its neighbours in $G^{\mathscr{R}}(P)$.
@@
For a partial solution $\mathscr{R}$ which does not completely cover $P$, a simplicial vertex always exists in $G^{\mathscr{R}}(P)$; and hence an unambiguous square given $\mathscr{R}$ always exists.
It is interesting to note that for a given simplicial node $p$ in $G^{\mathscr{R}}(P)$, there can be multiple maximal squares that cover $p$ and all its neighbours in $G^{\mathscr{R}}(P)$; they cover up the exact same set of uncovered blocks, but they may overlap differently with rec-packs in $\mathscr{R}$. Umabiguous squares are essentially equivalent to essential squares defined in the works of Bar-Yehuda and Ben-Chanoch [6].
@@
If $\mathscr{R}$ is a partial solution for an orthogonal polygon $P$ and $S$ is an unambiguous square given $\mathscr{R}$, then $\mathscr{R}\cup\{S\}$ is also a partial solution.
Proof 4.12.
Let $p$ be a simplicial node in $G^{\mathscr{R}}(P)$ such that $S$ covers $p$ and its neighbours in $G^{\mathscr{R}}(P)$. Since $\mathscr{R}$ is a partial solution, there exists a set of rec-packs $\mathscr{R}^{\prime}$ which is a minimum cover for $P$, satisfying $\mathscr{R}\subseteq\mathscr{R}^{\prime}$. Let $S^{\prime}$ be some square in $\mathsf{ext}(R^{\prime})$ that covers $p$.
Note that $S$ covers all such blocks in $G^{\mathscr{R}}(P)$ that $S^{\prime}$ covers. $(\mathsf{ext}(\mathscr{R}^{\prime})\setminus\{S^{\prime}\})\cup\{S\}$ therefore also forms a minimum cover of $P$. This implies that $\mathsf{ext}(\mathscr{R})\cup\{S\}$ as well as $\mathscr{R}\cup\{S\}$ are partial solutions.
Although we know that an unambiguous square given a partial solution $\mathscr{R}$ always exist, it is not trivial to algorithmically find one. The first step with which we achieve this is to get a polynomial-sized set of squares, which has at least one unambiguous square. The next result formally discusses this.
@@
Let $\mathscr{R}$ be a partial solution of an orthogonal polygon $P$ not completely covering $P$. We define $C_{x}$ and $C_{y}$ as follows.
@@
[…]
There exists an unambiguous square $S$ given $\mathscr{R}$ which has one corner in $C_{x}\cup C_{y}$.
Proof 4.14.
Notice that Proposition 2.16 guarantees the existance of at least one simplicial node in $G^{\mathscr{R}}(P)$. In fact, there can be multiple simplicial nodes. Let $V^{\prime}$ be the simplicial nodes with the largest number of neighbours in $G^{\mathscr{R}}(P)$ (highest degree). Among them, consider the topmost simplicial nodes in $V^{\prime}$, and let $p$ be the leftmost among them. Therefore, $p$ is the simplicial node in $G^{\mathscr{R}}(P)$ with the largest number of neighbours, and among the topmost of such simplicial nodes it is the leftmost.
Let $A$ denote the closed neighbourhood of $p$ in $G^{\mathscr{R}}(P)$. By Definition 2.14, $A$ must induce a clique in $G^{\mathscr{R}}(P)$, and hence in $G(P)$. Therefore, there must be a square that covers $A$ completely (Proposition 1.2). Let $S_{0}$ be a maximal square that covers $A$ completely, such that the position of the top-left corner of $S_{0}$ is leftmost among the topmost of all such squares covering $A$.
$S_{0}$ is maximal, so either the horizontal sides of $S_{0}$ individually overlap with two horizontal polygon edges, or the vertical sides of $S_{0}$ individually overlap with two vertical polygon edges (Lemma 4.5). We will assume that the horizontal sides of $S_{0}$ overlap with two horizontal polygon edges; the other case permits a symmetric argument.
We consider the following two cases.
@@
We first show that, in such a case, $G^{\mathscr{R}}(P)$ is a collection of isolated vertices.
Claim 1.
If $S_{0}$ is a $1\times 1$ square, then $G^{\mathscr{R}}(P)$ is a collection of isolated vertices.
Proof 4.15 (Proof of claim).
For the sake of contradiction, assume that $G^{\mathscr{R}}(P)$ has some arc. Consider the largest connected $G^{\star}$ of $G^{\mathscr{R}}(P)$. $G^{\star}$ induces a chordal graph, and hence must have a simplicial node, that has degree at least $1$. So the highest degree simplicial node in $G^{\mathscr{R}}(P)$ cannot be of degree 1 — contradiction.
Now, $p$ is the leftmost among the topmost of such simplicial nodes. Therefore, the immediate left block $p^{\prime}$ of $p$, is either (i) covered by a rec-pack in $\mathscr{R}$ or (ii) is outside the polygon.
If $p^{\prime}$ is covered by a rec-pack in $\mathscr{R}$, but $p$ is not, then the left side of $S_{0}$ is lies on the right side of some rec-pack $R$ in $\mathscr{R}$. Otherwise, if $p^{\prime}$ is outside the polygon, then the left side of $S$ lies on some polygon edge. Therefore, the top-left corner of $S_{0}$ must share the x-coordinate of either a corner of a rec-pack in $\mathscr{R}$ or a polygon vertex. Further, the top-left corner of $S_{0}$ shares a y-coordinate of a polygon vertex as the horizontal sides of $S_{0}$ overlap with polygon edges. This means that the top-left vertex of $S_{0}$ is in $C_{y}$.
@@
Recall that we assumed the horizontal sides of $S_{0}$ to overlap with polygon edges. If the left (resp. right) side of $S_{0}$ shares x-coordinate of some polygon edge or a corner of some rec-pack in $\mathscr{R}$, then the top-left (resp. top-right) corner of $S_{0}$ is in $C_{y}$; then we are done.
So, It suffices to assume that neither the left side nor the right side of $S_{0}$ share the x-coordinate with a polygon edge or with a corner of some rec-pack in $\mathscr{R}$. Let $S_{l}$ and $S_{r}$ be squares congruent to $S_{0}$, that are shifted one unit to the left and right respectively (Figure 6(a)).
Claim 2.
$S_{l}$ and $S_{r}$ are maximal valid squares.
Proof 4.16.
$S_{l}$ and $S_{r}$ are valid as the left and right sides of $S_{0}$ do not overlap with any polygon edge. Moreover, the horizontal sides of $S_{l}$ and $S_{r}$ overlap with polygon edges, the same edges with which the horizontal sides of $S_{0}$ overlap. So, $S_{l}$ and $S_{r}$ are maximal as well.
We further consider three more subcases.
  • Case II(a): All blocks inside $S_{0}$ along its right boundary, are also covered by some rec-pack in $\mathscr{R}$. Therefore, $S_{l}$ covers the same set of uncovered blocks as $S_{0}$. Moreover, the top-left corner of $S_{l}$ lies to the left of of the top-left corner of $S_{0}$. This contradicts the definition of $S_{0}$, and hence this subcase never arises.
  • Case II(b): All blocks inside $S_{0}$ along its left boundary, are also covered by some rec-pack in $\mathscr{R}$. Consider the following process: we keep moving $S_{0}$ horizontally to the right until one of the following happens:
    1. 1.
      it is obstructed by a polygon edge to its right.
    2. 2.
      moving it any further uncovers some previously covered block.
    3. 3.
      moving it any further will make one of its horizontal sides lose overlap with a polygon edge.
    Let the final square obtained be $S^{\prime}$. Due to the second condition, $S^{\prime}$ and $S_{0}$ cover the same set of uncovered blocks. Notice that since the process terminated with $S^{\prime}$, one of the three above conditions must hold for $S^{\prime}$. If further right movement is restricted due to a polygon edge, the right side of $S^{\prime}$ overlaps with a polygon edge. If the right movement is restricted as a covered block becomes uncovered, then the left side of $S^{\prime}$ overlaps with the right-side of some rec-pack of $\mathscr{R}$. Finally, if moving it any further makes $S^{\prime}$ lose overlap with a horizontal polygon edge, then there is a vertex of $P$, that coincides with the top-left corner or the bottom-left corner of $S^{\prime}$.
    Therefore, in all cases, one of the vertical sides share an x-coordinate of either a polygon vertex, or a rec-pack in $\mathscr{R}$. Since the horizontal sides of $S_{0}$ overlap with polygon edges, there is a corner of the $S^{\prime}$ in $\mathcal{C}_{y}$.
@@
    (b) Blocks $p_{l},p_{ll},p_{r},p_{rr}$ given $S_{0}$ and rec-packs in $\mathscr{R}$